OpenStudy (emilyjones284):

MEDAL AND FAN! (: The given series has six terms. What is the sum of the terms of the series? 17 + 25 + 33 + . . . + 57

OpenStudy (anonymous):

Same as: (17 + 0) + (17 + 8) + (17 + 16) + . . . + (17 + 40) = (17 + 0) + (17 + (8)(1)) + (17 + (8)(2)) + . . . + (17 + (8)(5)) = (6)(17) + 8(1 + 2 + 3 + 4 + 5) = (6)(17) + 8(5)(6)/2

OpenStudy (anonymous):

All good now, @emilyjones284 ?

OpenStudy (anonymous):

btw, 1 + 2 + 3 + . . . + (n - 2) + (n - 1) + n = (n)(n + 1)/2

OpenStudy (anonymous):

3(17+57)

OpenStudy (anonymous):

Good luck in all of your studies and thx for the recognition! @emilyjones284

OpenStudy (emilyjones284):

thanks and no problem!(: